Systems Staff Meeting Notes 05/12/03

Systems Staff Meeting Notes
May 12, 2003

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
  1. Cynthia has obtained tickets for “Matrix 2”.
  2. TEAMS meeting last week (more meetings have been scheduled) – they are extending the current structure to all of USPS. The seven generic categories will apply to positions: entry practitioner, etc. For further information, check the HRS Website on the UF site. This is a pilot program – could crash and burn for better or worse. Flexible hiring – contains some aspects of the federal system – none of this has a top end. 16-50 bottom end range, maximum was removed. IT was the first area to be revamped. Bill recommends keeping up with the changes. Hiring freeze? No absolute freeze, no statewide edict. Exceptions can be made. Change jobs at your own risk. Remember: last hired, first fired. FSU – announcing layoffs in July. There is no state budget yet – special session is necessary. Uncertain times.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Dell has finally acknowledged that they have received our order for 5 PowerEdge servers. No estimated ship date is available at this time.
  1. Our anti-virus vendor released a warning about the Fizzer A mail/file sharing worm. Our current version of Trend should catch it.
  2. MSL 194L drops complete. They wanted to move the desks once the drops were started. Orange picnic fencing will direct traffic. The whole project is proceeding in free-form fashion. Xerox’s new location is unknown – maybe on the 4th or 5th floor. No notion of storage area yet – current plan is 250,000 volumes to Room 1 East. Plans change continuously; we need to react as quickly as we can to any and all requests. Bobbie met with Tim to figure out the wiring costs – general estimate $34,000 worth on the first floor: 144 pulls, 100 data drops, 44 phones, patch panels and closet. This seems like a very high estimate. Need to re-visit those figures. Areas will start closing down on May 15. The move is scheduled to start on May 19 (next Monday).
  3. WINS server is fine as of noon last Monday.
  4. Will has the Exchange recovery server built and is working on replication. Then it will be moved to MSL to backup Exchange, IE, IIS Web server and SQL.
  5. Moving Exchange to NT2R – cluster version – there’s a trick to resetting the security – Bill knew it – which help to fix things. Update can reset local security – sometimes a template will fix it.
Software Projects Status
  1. SysRequest Rollout – Cynthia spoke with Gary Cornwell – he’s going out of town. Cynthia will train Leilani, who will pass it on to Gary. No word from her as of yet. Bill instructs to keep going down the line, training everybody who can use the database. Need to work on the multiple workstation entry.
  2. Facilities – will allow order point, incoming field with notation. They are still technically in test mode, but they are using it. There is one more critical component – altering the invoices after they go out – then it should be ready to go live.
  3. Geoffrey, Winston and Hilton – trying auto-format for security on public workstations. They are getting closer to success. We’ll be ordering more disk drives (about two dozen) for the Info Commons.
  4. ALEPH support – Rich met last Friday with FCLA. They are trying to set up the UF ID into NOTIS record by the Summer B semester.
  5. Winston sent the first Patron Update records today.
  6. DLC – postcards are done. Aerial is waiting for a meeting with Joe Aufmuth (Docs) and Erich. Both are out this week. Sanborne is ongoing, there are still some questions.
  7. Geoffrey notes that DLC has updated their scanning software. If they have a problem, don’t get involved – it’s a driver issue.
Web Projects
  1. One drive on the Stat server is close to full. Reports won’t work. Suzy is attending to it.
  2. Webmaster position – interviews this week, it closes May 30. They want to hire two weeks after the position closes. Will ask Denise about the hiring freeze.
  3. Documents is back online. They somehow blew away both of their Web sites while backing them up on Friday. They do no know how it happened. We are still reloading from our own backups.
Student and Department Happenings
  1. New student – Nikia Wilson – started this morning.

Back to the Systems Homepage

Last updated May 30, 2003
by Debra Fetzer